Grade 11General Physics

A beam of light converges at a point P. Now a lens is placed in the path of the convergent beam 12cm from P. At what point does the beam converge if the lens is (a) a convex lens of focal length 20cm, and (b) a concave lens of focal length 16cm?

Jayant Kumar
12 Years agoGrade 11

1 Answer

Malhar Chakraborty
9 Years ago
Object distance, u = +12 cm (virtual object)
 
(a) Focal length of the convex lens, f = 20 cm
     Image distance = v
     Applying lens formula 
     1/v – 1/u = 1/f
     Therefore putting the values and solving the equation, we get v = 60/8 i.e. = 7.5 cm
     Hence, the image is formed 7.5 cm away from the lens, toward its right.
(b) Focal length of the concave lens, f = -16
     again image distance = v
     the again as previous, applying the lens formula, substituing the values of u and f, we get v= 48 cm.
     Hence, the image is formed 48 cm away from the lens, towards its right.